OpenStudy (anonymous):

plzzzz need help ODEs \((\cos x) (e^y -y) \frac{dy}{dx}=e^y\sin 2x \)

ganeshie8 (ganeshie8):

separate variables

OpenStudy (anonymous):

\( \frac{(e^y -y)}{e^y} dy =\frac {\sin 2x}{\cos x} dx \)

terenzreignz (terenzreignz):

Doesn't that look simpler? Here... let me get that for you ... \[\Large (1 - ye^{-y})dy = \frac{2\sin(x)\cos(x)}{\cos(x)}dx\] better? ^_^

OpenStudy (anonymous):

much better <3

OpenStudy (anonymous):

so nw i only need integration by part for LHS RHS is much simple right ?

terenzreignz (terenzreignz):

Of course :) So I take it you can do it from here?

OpenStudy (anonymous):

yes :D thanks all of u guys